# Embedding-based statistical inference on generative models

### v1

@@ Abstract
The recent cohort of publicly available generative models can produce human expert level content across a variety of topics and domains. Given a model in this cohort as a base model, methods such as parameter efficient fine-tuning, in-context learning, and constrained decoding have further increased generative capabilities and improved both computational and data efficiency. Entire collections of derivative models have emerged as a byproduct of these methods and each of these models has a set of associated covariates such as a score on a benchmark, an indicator for if the model has (or had) access to sensitive information, etc. that may or may not be available to the user. For some model-level covariates, it is possible to use “similar” models to predict an unknown covariate. In this paper we extend recent results related to embedding-based representations of generative models – the data kernel perspective space – to classical statistical inference settings. We demonstrate that using the perspective space as the basis of a notion of “similar” is effective for multiple model-level inference tasks.
corresponding author
@@ 1 Introduction
Generative models have recently met or surpassed human-level standards on benchmarks across a range of tasks (Nori et al., 2023; Katz et al., 2024; Dubey et al., 2024). While these claims warrant skepticism and further robustness evaluation (Ness et al., 2024), the impressive capabilities have created a competitive environment for training state-of-the-art models and have inspired the development of complementary methods to adapt models to particular use cases. For example, quantizing the weights of a neural model enables a trade-off of model precision with vRAM and disk space (Gholami et al., 2022) while methods such as Low Rank Adaptation (LoRA) (Hu et al., 2021) and prompt-tuning (Lester et al., 2021) enable compute- and data-efficient model adaptation. Other methods such as retrieval-augmented generation (Lewis et al., 2020), model merging (Matena & Raffel, 2022), constrained decoding (Hokamp & Liu, 2017), etc. (Dettmers et al., 2024; Edge et al., 2024), have similarly contributed to the rapid development of a large population of diverse and accessible models.
Each model in the population has an accompanying set of covariates – scores on benchmarks, training mixture proportions, model safety scores, probability of hallucination, etc. – that are a function of the model, the training set, the architecture, the retrieval database, or derivatives thereof. For a given model, the covariate of interest or the function to calculate it may not be available to the user. For example, a proprietary model may have been trained on copyrighted data and the indicator for whether or not the model has had access to the data is unknown. Methods for predicting model-level covariates are necessary in these settings, and others, to fully understand the behavior and properties of a model.
In this paper we extend recent theoretical results for embedding-based representations of generative models (Acharyya et al., 2024) to statistical inference settings. In particular, our results show that the embedding-based representations of a collection of models can be used for consistent inference for a wide class of inference problems. We demonstrate the effectiveness of the representations for three downstream inference tasks, including predicting the presence of sensitive information in a model’s training mixture and predicting a model’s safety. We include empirical investigations of performance sensitivity to hyperparameters required to induce the model representations.
@@ 1.1 Background & related work
Lastly, our work is a part of the relatively new literature on inference on generative models. For example, FlashHELM (Perlitz et al., 2023) uses the score on a subset of a benchmark such as HELM (Liang et al., 2022) to quickly identify where a model fits on a leaderboard. More recent work proposes scaling laws to predict the performance of base models on a suite of benchmarks as a function of training FLOPs (Ruan et al., 2024). The method proposed herein does not assume access to a scoring function nor does it assume access to a function of the model’s weights at inference time. Perhaps most importantly, our setting and method can be applied to general model-level prediction scenarios.
@@ 2 The Data Kernel Perspective Space
Following Acharyya et al. (2024), we consider a generative model $f$ with weights $W$ and decoding function (temperature, maximum response length, etc.) $\delta$ to be a random map from a query space $\mathscr{Q}$ to a response space $\mathcal{X}$. We let $f_{1},\ldots,f_{n}$ be $n$ such random maps for a shared $\mathscr{Q}$ and $\mathcal{X}$. For example, $f_{i}$ may be the neural model parameterized by the fixed context augmentation $a_{i}$, the retrieval vector database $V_{i}$, the fine-tuned model with weights $W+\Delta W_{i}$ or the model with decoding function $\delta_{i}$. Given a collection of queries $Q=\{q_{1},\ldots,q_{m}\}$ we let $F_{ij}$ denote the distribution on the response space for fixed $f_{i}$ and $q_{j}$ and assume model responses $f_{i}(q_{j})_{1},\ldots,f_{i}(q_{j})_{r}=x_{ij1},\ldots,x_{ijr}$ are $i.i.d.$ realizations from $F_{ij}$.
Our goal is to induce a representation of the neural models that captures information relevant to model-level statistical inference. We abuse notation by letting $x$ represent a realization of a random variable and the random variable itself and include necessary context when there is potential for confusion.
@@
We let $g:\mathcal{X}\to\mathbb{R}^{p}$ be an embedding function that maps a response to a $p$-dimensional real-valued vector and, for notational convenience, let $\bar{x}_{ij}=\frac{1}{r}\sum_{k=1}^{r}g(x_{ijk})$ be the average of the embedded responses. We use $\bar{X}_{i}\in\mathbb{R}^{m\times p}$ to denote the matrix where the $j$th row is $\bar{x}_{ij}$ and view this matrix as a representation of model $f_{i}$ with respect to $Q$ and $g$.
We define $D$ as the $n\times n$ pairwise distance matrix with entries
$D_{ii^{\prime}}=\frac{1}{m}\big{|}\big{|}\bar{X}_{i}-\bar{X}_{i^{\prime}}\big{|}\big{|}_{F}.$ (1)
@@
Each entry $D_{ii^{\prime}}$ is the average difference between the average embedded response between model $f_{i}$ and model $f_{i^{\prime}}$ and captures the difference in average model behavior with respect to $Q$ and $g$. With the form described by Eq. (1), $D$ is a Euclidean distance matrix and the multidimensional scaling (MDS) of $D$ yields $d$-dimensional Euclidean representations of the matrices $\bar{X}_{i}$ (Torgerson, 1952) and, thus, Euclidean representations of the models $f_{1},\ldots,f_{n}$ with respect to $Q$ and $g$. Letting $\widehat{\psi}:=\text{MDS}(D)$, we refer to the $i$th row of $\widehat{\psi}_{i}\in\mathbb{R}^{d}$ as the perspective of model $f_{i}$ and the entire configuration $\widehat{\psi}$ as the data kernel perspective space (DKPS).
2.1 Analytical properties of the data kernel perspective space
While the perspectives are Euclidean objects, it is not possible to comment on their properties without imposing constraints on the $F_{ij}$. To facilitate analysis we assume each query is from the query distribution $G$ with support $\mathscr{Q}$ and let $\mu_{i}\in\mathbb{R}^{m\times p}$ denote the matrix whose $i^{th}$ row is $\mathbb{E}_{x\sim F_{ij}}\left[g(x)\right]$. We let $\Delta$ be such that $\Delta_{ii^{\prime}}=\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$. For settings where $m$ grows and $q_{j}\overset{iid}{\sim}G$, we refer to $\mu_{i}$ as $\mu_{i}(G)$ and let $\Delta^{*}(G)$ be such that $\Delta^{*}_{ii^{\prime}}(G)=\lim_{m\to\infty}\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$. Finally, we assume $g$ is bounded. [^1] Boundedness is typically satisfied in practice for well-defined $g$: for language models, an element of $\mathcal{X}$ is a finite sequence from a finite vocabulary; for text-to-image models $\mathcal{X}=\{0,\ldots,255\}^{3}$ is finite.
With $q_{1},\ldots,q_{m}\overset{iid}{\sim}G$, and $x_{ij1},\ldots,x_{ijr}\overset{iid}{\sim}F_{ij}$, we have
$D_{ii^{\prime}}=\frac{1}{m}\left\lVert\bar{X}_{i}-\bar{X}_{i^{\prime}}\right\rVert\to\Delta^{*}_{ii^{\prime}}(G)$ (2)
@@ 2.1 Analytical properties of the data kernel perspective space
with high probability as $m,r\to\infty$ for all $(i,i^{\prime})\in\{1,\ldots,n\}\times\{1,\ldots,n\}$. Under technical assumptions described in Acharyya et al. (2024), there exists $\psi(G):=\text{MDS}(\Delta^{*}(G))$ such that $\widehat{\psi}\to\psi(G)$. The configuration $\psi(G)$ captures the true geometry of the mean descrepancies of the model responses with respect to queries from $G$. We drop the argument of $\psi$ when the distribution of the queries is clear. In a more general setting where the models are assumed to be $i.i.d.$ realizations from a model distribution $F_{model}$, the distance matrix $D$ converges to $\Delta^{*}$ and, under technical assumptions, there exists $\psi$ such that $\widehat{\psi}\to\psi$ as $m,r\to\infty$, for all $n$ (Acharyya et al., 2024).
In settings where the query space contains only a single element or where each $F_{ij}$ is a point mass on $x_{ij}$, $\Delta_{ii^{\prime}}$ is exactly the maximum mean discrepancy (Gretton et al., 2012) of the distributions $F_{ij}$ and $F_{i^{\prime}j}$ under a linear map applied to $g(x)$. For settings where $|\mathscr{Q}|>1$ and the $F_{ij}$ are not all point masses, the analysis is more complicated since both $F_{ij}$ and $\bar{x}_{ij}$ are random variables.
2.2 Statistical inference in the data kernel perspective space
@@
Consider the classical statistical learning problem (Hastie et al., 2009, Chapter 2): Given training data $\mathcal{T}_{n}=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ assumed to be $i.i.d.$ realizations from the joint distribution $P_{XY}$, select the decision function $h:\mathcal{X}\to\mathcal{Y}$ that minimizes the expected value of a loss function $\ell$ with respect to $P_{XY}$ within a class of decision functions $\mathcal{H}$ for a test observation $X$ assumed to be an independent realization from the marginal distribution $P_{X}$. Or, with $\mathcal{R}_{\ell}(P_{XY},h):=\mathbb{E}_{P_{XY}}\left[\ell(h(X),y)\right]$, select $h$ such that
$h\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\mathcal{R}_{\ell}(P_{XY},h).$
We let $\mathcal{R}_{\ell}^{*}(P_{XY},\mathcal{H})$ denote the expected loss of $h$ in the argmin and let $\mathcal{R}^{*}_{\ell}(P_{XY})$ denote the minimum $\mathcal{R}_{\ell}^{*}(P_{XY},\mathcal{H})$ over all $\mathcal{H}$.
@@
The joint distribution $P_{XY}$ is often unavailable and the decision function is selected based on $\mathcal{T}_{n}$. We let $h(\;\cdot\;;\mathcal{T}_{n})$ denote such a decision function and say the sequence of decision functions $\left(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n})\right)$ is consistent for $P_{XY}$ with respect to $\mathcal{H}$ if
$Pr\big{(}\;\big{|}\mathcal{R}_{\ell}(P_{XY},h(\;\cdot\;;\mathcal{T}_{n}))-\mathcal{R}^{*}_{\ell}(P_{XY},\mathcal{H})\big{|}>\epsilon\big{)}\to 0$
@@
In our setting we observe $\widehat{\mathcal{T}}_{n}=\{(\widehat{\psi}_{1},y_{1}),\ldots,(\widehat{\psi}_{n},y_{n})\}$ and the test observation $\widehat{\psi}_{n+1}$ where $\widehat{\psi}_{i}$ is an errorful observation of the true-but-unknown $\psi_{i}$ and $y_{i}$ is the model-level covariate corresponding to $f_{i}$. Two important extensions of the convergence results described in (Acharyya et al., 2024) are the following theorems:
@@
as $m,r\to\infty$;
and
@@
Under technical assumptions described in Appendix A, if $(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ then $(h(\;\cdot\;;\widehat{\mathcal{T}}_{1}),\ldots,h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ as $n,m,r\to\infty$.
@@
These results do not say that with enough models, queries, and responses that inference using the $\widehat{\psi}$ is “optimal”. Instead, Theorem 1 says that with enough queries and responses for each query that the performance of a decision function that uses $\widehat{\mathcal{T}}_{n}$ will be “close” to the performance of the decision function that uses $\mathcal{T}_{n}$ – regardless of how far away $\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))$ is from $\mathcal{R}_{\ell}^{*}(P_{\psi Y},\mathcal{H})$.
Similarly, Theorem 2 says that if there is a sequence of decision functions known to be consistent with respect to $\mathcal{H}$ when using $\mathcal{T}_{n}$, then the same sequence of decision functions is consistent with respect to $\mathcal{H}$ when using $\widehat{\mathcal{T}}_{n}$ where $m$ and $r$ are allowed to grow. The result does not provide instructions on how to construct such a sequence. Nor does it provide insight as to how close $\mathcal{R}_{\ell}^{*}(P_{\psi Y},\mathcal{H})$ is to $\mathcal{R}_{\ell}^{*}(P_{\psi Y})$. Most importantly – and most unique to our setting – Theorem 2 does not provide guidance for how to select a query distribution $G$ such that $|\mathcal{R}_{\ell}^{*}(P_{fY})-\mathcal{R}_{\ell}^{*}(P_{\psi Y})|$ is minimal.
While we do not provide theoretical details as to which situations optimal model performance is possible, if $y_{i}$ can be described as a function of the average of model responses for a set of queries from an optimal $G^{*}$ then $\psi(G^{*})$ will be such that $\mathcal{R}_{\ell}^{*}(P_{fY})=\mathcal{R}_{\ell}^{*}(P_{\psi Y})$ (Devroye et al., 2013, Chapter 32). Conversely, it is unclear how effective decision functions constructed with $\psi(G^{*})$ for model-level covariates that cannot be described as a function of the average response will behave. Further, it is unclear how “close” a query distribution $G$ needs to be to $G^{*}$ to induce an informative representation of the models with respect to $y_{i}$. We demonstrate a potential trade-off between the proximity of $G$ to $G^{*}$ and the number of queries necessary to tease out model differences in one of the experimental settings below.
Lastly, we note that our setting, methodology, and theorems are general to classes of generative models in which there exists a well-defined embedding function from the output space to $\mathbb{R}^{p}$, though our experimental settings focus entirely on collections of language models.
2.3 An illustrative example – “Was RA Fisher great?”

### v3

# Statistical inference on black-box generative models in the data kernel perspective space

@@ Abstract
Generative models are capable of producing human-expert level content across a variety of topics and domains. As the impact of generative models grows, it is necessary to develop statistical methods to understand collections of available models. These methods are particularly important in settings where the user may not have access to information related to a model’s pre-training data, weights, or other relevant model-level covariates. In this paper we extend recent results on representations of black-box generative models to model-level statistical inference tasks. We demonstrate that the model-level representations are effective for multiple inference tasks.
Statistical inference on black-box generative models in the data kernel perspective space
Hayden Helm Helivan hayden@helivan.io                        Aranyak Acharyya Johns Hopkins University                        Youngser Park Johns Hopkins University
Brandon Duderstadt Nomic AI                        Carey E. Priebe Johns Hopkins University
@@ 1 Introduction
Generative models have recently met or surpassed human-level standards on benchmarks across a range of tasks (Nori et al., 2023; Katz et al., 2024; Dubey et al., 2024). While these claims warrant skepticism and further robustness evaluation (Ness et al., 2024), the impressive capabilities have created a competitive environment for training state-of-the-art models and have inspired the development of complementary methods to adapt models to particular use cases. For example, quantizing the weights of a neural model enables a trade-off of model precision with vRAM and disk space (Gholami et al., 2022) while methods such as Low Rank Adaptation (LoRA) (Hu et al., 2021) and prompt-tuning (Lester et al., 2021) enable compute- and data-efficient model adaptation. Other methods such as retrieval-augmented generation (Lewis et al., 2020), model merging (Matena and Raffel, 2022), constrained decoding (Hokamp and Liu, 2017), etc. (Dettmers et al., 2024; Edge et al., 2024), have similarly contributed to the rapid development of a large population of diverse and accessible models.
Each model in the population has an accompanying set of covariates – scores on benchmarks, training mixture proportions, model safety scores, etc. – that are a function of the model, the training set, the architecture, the retrieval database, or derivatives thereof. For a given model, the covariate of interest may not be available to the user or it may be too expensive to calculate. For example, it may not be known if a proprietary model has been trained on copyrighted data. Or, it may be too resource intensive to directly evaluate the toxicity of every model. Methods for predicting model-level covariates are necessary in these settings, and others, to fully understand the behavior and properties of a model.
In this paper we extend recent theoretical results for vector representations of black-box generative models (Acharyya et al., 2024) to model-level statistical inference settings. Our results show that the embeddings of the responses from a collection of generative models can be used for consistent inference for a wide class of inference problems. We demonstrate the effectiveness of the representations for three downstream inference tasks, including predicting the presence of sensitive information in a model’s training mixture and predicting a model’s safety. We include empirical investigations of performance sensitivity to hyperparameters required to generate the model-level vector representations.
Contribution. Our contribution is the theoretical and empirical validation for using vector representations of black-box generative models for model-level inference tasks. The representations that we study herein are based on the embeddings of their responses. The theoretical results apply to any generic, well-behaved embedding function. Given the representations, any vector-based method can be used for inference on the models.
@@ 1.1 Background & related work
Lastly, our work is a part of the relatively new literature on inference on generative models. For example, FlashHELM (Perlitz et al., 2023) uses the score on a subset of a benchmark such as HELM (Liang et al., 2022) to quickly identify where a model fits on a leaderboard. More recent work proposes scaling laws to predict the performance of base models on a suite of benchmarks as a function of training FLOPs (Ruan et al., 2024). The method proposed herein does not assume access to a scoring function nor does it assume access to a function of the model’s weights at inference time. Perhaps most importantly, our setting and method can be applied to general model-level prediction.
@@ 2 The Data Kernel Perspective Space
Before we present our results related to statistical inference on black-box generative models, we first describe how to obtain finite-dimensional vector representations of the models. The particular method for obtaining vector representations for generative models that we study herein has previously been discussed in Helm et al. (2024) and Acharyya et al. (2024).
Let $f:\mathcal{Q}\to\mathcal{X}$ be a generative model from a query space $\mathcal{Q}$ to an output space $\mathcal{X}$. Given $q\in\mathcal{Q}$, we assume that $f(q)$ produces a random element in $\mathcal{X}$. In our setting there are $n$ models $f_{1},\ldots f_{n}$ and $m$ queries $q_{1},\ldots,q_{m}$ with $q_{j}\in\mathcal{Q}$. We assume that each model responds to every query $r$ independent times and let $f_{i}(q_{j})_{k}$ be the $k$-th response to $q_{j}$ from $f_{i}$.
@@
We let $g:\mathcal{X}\to\mathbb{R}^{p}$ be an embedding function that maps a model response to a $p$-dimensional real-valued vector. Further, we let $x_{ijk}=g(f_{i}(q_{j})_{k})$ denote the embedding of $f_{i}(q_{j})_{k}$ and $F_{ij}$ denote the distribution of $x_{ijk}$ on $\mathbb{R}^{p}$. That is, $x_{ij1},\dots x_{ijr}\overset{iid}{\sim}F_{ij}$.
We let $\bar{x}_{ij}=\frac{1}{r}\sum_{k=1}^{r}x_{ijk}$ be the empirical average embedded response of $f_{i}$ to $q_{j}$ and $\bar{X}_{i}\in\mathbb{R}^{m\times p}$ denote the matrix whose $j$-th row is $\bar{x}_{ij}$. We view $\bar{X}_{i}$ as a matrix representation of model $f_{i}$ with respect to $\{q_{1},\dots,q_{m}\}$ and $g$. We define $D$ as the $n\times n$ pairwise distance matrix with entries
$D_{ii^{\prime}}=\frac{1}{m}\big{|}\big{|}\bar{X}_{i}-\bar{X}_{i^{\prime}}\big{|}\big{|}_{F}$ (1)
and note that $D_{ii^{\prime}}$ captures the difference in empirical average model behavior between $f_{i}$ and $f_{i^{\prime}}$ with respect to $\{q_{1},\ldots,q_{m}\}$ and $g$.
@@
With the form described by Eq. (1), $D$ is a rescaled Euclidean distance matrix. Hence, multidimensional scaling (MDS) of $D$ yields $d$-dimensional Euclidean representations of the matrices $\bar{X}_{i}$ (and thereby of the models $f_{i}$) with respect to $\{q_{1},\ldots,q_{m}\}$ and $g$ (Torgerson, 1952). Letting $\widehat{\psi}:=\text{MDS}(D)\in\mathbb{R}^{n\times d}$, we refer to $\widehat{\psi}$ as the data kernel perspective space (DKPS). We emphasize that the $i$-th row of the DKPS – $\widehat{\psi}_{i}$ – is a $d$-dimensional vector representation of the generative model $f_{i}$.
2.1 Analytical properties of the DKPS
While the DKPS representations of the models are Euclidean objects, it is not possible to comment on their properties without imposing constraints on the $F_{ij}$. We let $\mu_{i}\in\mathbb{R}^{m\times p}$ be the population counterpart of $\bar{X}_{i}$ whose $j$-th row is $\mathbb{E}_{x\sim F_{ij}}(x)$. Similarly, we let $\Delta$ be the population counterpart of $D$ whose $(i,i^{\prime})$-th element is $\Delta_{ii^{\prime}}=\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$. We assume $\Delta^{*}_{ii^{\prime}}=\lim\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$ for every pair $(i,i^{\prime})$ as $m,r\to\infty$ exists and that $g$ is bounded.
In our setting, under appropriate assumptions described in (Acharyya et al., 2024),
$D_{ii^{\prime}}=\frac{1}{m}\left\lVert\bar{X}_{i}-\bar{X}_{i^{\prime}}\right\rVert\to\Delta_{ii^{\prime}}^{*}$ (2)
@@ 2.1 Analytical properties of the DKPS
with high probability as $m,r\to\infty$ for all $(i,i^{\prime})\in\{1,\ldots,n\}\times\{1,\ldots,n\}$. Further, there exists $\psi:=\text{MDS}(\Delta^{*})\in\mathbb{R}^{m\times d}$ such that $\widehat{\psi}\to\psi$. That is, the configuration $\psi$ is the population counterpart of $\widehat{\psi}$. $\psi$ captures the true geometry of the mean discrepancies of the model responses with respect to the queries $\{q_{1},\ldots,q_{m}\}$. Importantly, in settings where the models are not fixed and are instead assumed to be $i.i.d.$ realizations from a model distribution $P_{model}$, the distance matrix $D$ converges to $\Delta^{*}$ and, under technical assumptions, there exists $\psi$ such that $\widehat{\psi}\to\psi$ as $m,r\to\infty$ for all $n$ (Acharyya et al., 2024).
2.2 Statistical inference in the DKPS
@@
We now extend the consistency of $\widehat{\psi}$ to $\psi$ to model-level inference. Consider the classical statistical learning problem (Hastie et al., 2009, Chapter 2) in the context of a collection of generative models: given training data $\mathcal{T}_{n}=\{(f_{1},y_{1}),\ldots,(f_{n},y_{n})\}$ assumed to be $i.i.d.$ realizations from the joint distribution $P_{fY}$, choose the decision function $h:\mathcal{F}\to\mathcal{Y}$ that minimizes the expected value of a loss function $\ell$ with respect to $P_{fY}$ within a class of decision functions $\mathcal{H}$ for a test observation $f$ assumed to be an independent realization from the marginal distribution $P_{f}$. Or, with $\mathcal{R}_{\ell}(P_{fY},h):=\mathbb{E}_{P_{fY}}\left[\ell(h(f),y)\right]$, select $h^{*}$ such that
$h^{*}\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\mathcal{R}_{\ell}(P_{fY},h).$
We let $\mathcal{R}_{\ell}^{*}(P_{fY},\mathcal{H})=\mathcal{R}_{\ell}(P_{fY},h^{*})$ denote the expected loss (or “risk") of $h^{*}$.
@@
The joint distribution $P_{fY}$ is often unavailable and the decision function is selected based on the training data. We let $h(\;\cdot\;;\mathcal{T}_{n})$ denote such a decision function and say the sequence of decision functions $\left(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n})\right)$ is consistent for $P_{fY}$ with respect to $\mathcal{H}$ if
$Pr\big{(}\;\big{|}\mathcal{R}_{\ell}(P_{fY},h(\;\cdot\;;\mathcal{T}_{n}))-\mathcal{R}^{*}_{\ell}(P_{fY},\mathcal{H})\big{|}>\epsilon\big{)}\to 0$
@@
In the black-box setting we do not have direct access to useful representations of the models. Instead, we must query each model repeatedly and use the representations $\{\psi_{1},\ldots,\psi_{n}\}$ as proxies for the models $\{f_{1},\ldots,f_{n}\}$. Hence, $\mathcal{T}_{n}=\{(\psi_{1},y_{1}),\ldots,(\psi_{n},y_{n})\}$ where $(\psi_{i},y_{i})\sim^{iid}P_{\psi Y}$ and our goal is to choose a decision function $h^{*}$ which minimizes the average loss $\mathcal{R}_{\ell}(P_{\psi Y},h):=\mathbb{E}_{P_{\psi Y}}\left[\ell(h(\psi),y)\right]$ within an appropriately defined $\mathcal{H}$. Note that the true $\{\psi_{1},\ldots,\psi_{n}\}$ are not known a priori and must be estimated via $\{\widehat{\psi}_{1},\ldots,\widehat{\psi}_{n}\}$. We let $\widehat{\mathcal{T}}_{n}$ be the training data where $\psi_{i}$ is replaced with $\widehat{\psi}_{i}$.
Two important extensions of the DKPS consistency results in the context of the classical statistical learning problem are the following theorems:
@@
as $m,r\to\infty$, for every $n$.
That is, the risk of a decision function based on $\widehat{\mathcal{T}}_{n}=\{(\widehat{\psi}_{1},y_{1}),\ldots,(\widehat{\psi}_{n},y_{n})\}$ converges to the risk of the decision function based on the true-but-unknown $\mathcal{T}_{n}$. For situations where $\{\psi_{1},\ldots,\psi_{n}\}$ are good proxies, Theorem 1 states that increasing the number of queries and/or number of replicates per query will improve the performance of decision functions trained on $\widehat{\mathcal{T}}_{n}$.
@@
Under technical assumptions described in Appendix A, if $(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$, then $(h(\;\cdot\;;\widehat{\mathcal{T}}_{1}),\ldots,h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ as $n,m,r\to\infty$.
Theorem 2 states that if the sequence of decision functions learned from $\mathcal{T}_{n}$ is consistent then the sequence of decision functions learned from the analogous $\widehat{\mathcal{T}}_{n}$ is also consistent. This result suggests that inference performance should improve with more training data. While a logical follow-on to Theorem 1 and a well-understood principle in classical statistical learning settings, the consistency of the DKPS in the setting where the number of models grows is a non-trivial extension of the fixed $n$ case.
@@
2.2.1 Non-standard considerations
Given that our analysis is based on estimates $\{\widehat{\psi}_{1},\ldots,\widehat{\psi}_{n}\}$ of proxies $\{\psi_{1},\ldots,\psi_{n}\}$ of the objects of interest $\{f_{1},\ldots,f_{n}\}$, it is important to note some non-standard theoretical considerations. We let $\mathcal{R}^{*}_{\ell}(P_{fY}):=\inf_{\{h:\mathcal{F}\to\mathcal{Y}\}}\mathcal{R}_{\ell}(P_{fY},h)$ be the Bayes optimal risk of $P_{fY}$ and $P_{query}$ be a probability distribution on queries.
For example, while our theoretical results provide insights as to how performance will be affected by increasing $n,m,$ and $r$, our results do not comment on the magnitude of $|\mathcal{R}^{*}_{\ell}(P_{\psi Y})-\mathcal{R}^{*}_{\ell}(P_{fY})|$, which measures how good of a proxy $\{\psi_{1},\ldots,\psi_{n}\}$ is for $\{f_{1},\ldots,f_{n}\}$. In particular, if $|\mathcal{R}^{*}_{\ell}(P_{\psi Y})-\mathcal{R}^{*}_{\ell}(P_{fY})|=0$ then the $\psi_{i}$ are lossless proxies of the $f_{i}$ with respect to $y$ (Devroye et al., 2013, Chapter 32). We expect it to be challenging to understand $|\mathcal{R}^{*}_{\ell}(P_{\psi Y})-\mathcal{R}^{*}_{\ell}(P_{fY})|$ in a general setting.
Similarly, let $q_{1},\ldots,q_{m}\sim^{iid}P_{query}$ and $q^{\prime}_{1},\ldots,q^{\prime}_{m}\sim^{iid}P^{\prime}_{query}$ with $P_{query}\neq P^{\prime}_{query}$. The proxies of the models induced by these two query sets – $\{\psi_{1},\ldots,\psi_{n}\}$ and $\{\psi_{1}^{\prime},\ldots,\psi_{n}^{\prime}\}$, respectively – will be of different quality. Without an initial understanding of the models and their relationship with the covariates, we expect that it will be impossible to know which query distribution is preferred a priori. We highlight the importance of $P_{query}$ in an experimental setting below.
Lastly, we note that representations of the models that we analyze herein capture the intrinsic geometry of the mean discrepancies of the models for the queries $\{q_{1},\ldots,q_{n}\}$. For covariates that cannot be described as a function of the mean discrepancies of the models, it is unclear how to interpret Theorems 1 and 2. We discuss potential alternatives to the representations that we study in the limitations section.
2.3 An illustrative example – “Was RA Fisher great?"